A sound source emits sounds of frequency 210 Hz that travel through still air at 340 m/s. The listener moves at 70 m/s relative to still air toward the stationary source. What is the wavelength of the sound between the source and the listener? (not 1.286 or 264.39)
Answer #1

As the listener moves towards the source the wavelength remains same


wavelength lambda = v/f

v = speed of sound = 340 m/s

f = frequency of sound

lambda = 340/210 = 1.619 m <<<-------ANSWER